OpenStudy (anonymous):

show that 50x/(20-x) = 1000/(20-x)-50

OpenStudy (anonymous):

\[\frac{ 50x }{20-x }=\frac{ 1000 }{ 20-x }-50\]

OpenStudy (anonymous):

450x + 1000 = 500x 1000 = 50x 20 = x there must be 20 in sales C = 450(20) + 1000 = 10,000 R = 500(20) = 10,000 any sales above 20 means that revenue will be greater than cost, and the business will make a profit

OpenStudy (anonymous):

what?